A work force of 16 men is required to complete an assignment in 10 days. By how much should the work force be reduced in order to complete the assignment in 40 days?

Grade:7

1 Answers

Shivansh
5 years ago
16*10 = n*40
where n is the number of men required
hence n is 4
the workforce should be reduced by 16/4 = 4 times